# Imaginary Gauge-steerable Edge Modes In Non-Hermitian Aubry-André-Harper Model

### v1

@@ VI Conclusion
We have studied a quasiperiodic non-Hermitian AAH chain with a zero-mean, spatially fluctuating imaginary gauge field. Under OBC, the model was similarity-equivalent to the Hermitian AAH chain, so the OBC spectrum was preserved while the right/left eigenstates acquired a realization-dependent random-walk envelope. In gaps that host isolated boundary modes in the Hermitian reference problem, we found two in-gap modes with distinct responses: one remained boundary pinned, whereas the other was gauge steerable, with its probability maximum shifted by changing only the gauge realization without moving its in-gap eigenvalue. We also showed that the biorthogonal weight controlling the first-order response to a local on-site perturbation reduced exactly to the site-resolved density of the corresponding Hermitian eigenstate, which fixed the optimal gain site independently of the gauge realization and enabled a weak local-gain protocol to prepare the steerable mode from a bulk initial state. Our work uncovers an exactly tractable route to position-tunable in-gap localization in non-Hermitian quasiperiodic lattices and opens a new avenue to analytically explore erratic-gauge boundary physics with clear experimental feasibility.
@@ Acknowledgements.
X.L.Z. thanks Zhi Li and Peng Zou for helpful discussions. This work was supported by the Joint Fund of Natural Science Foundation of Shandong Province (Grant No. ZR2024LLZ004), the National Natural Science Foundation of China(Grant No. 12005110), and the Natural Science Foundation of Shandong Province (Grant No. ZR2020QA078, No. ZR2023QB065, No. ZR2023MD064).

### v2

@@ VI Conclusion
We have investigated a quasiperiodic non-Hermitian AAH chain with a zero-mean, spatially fluctuating imaginary gauge field under OBC. An exact nonunitary gauge transformation mapped the system to its Hermitian AAH counterpart: the spectrum and Lyapunov exponents were unchanged, while the right/left eigenstates were reshaped by the gauge-induced envelope. In a parameter region with spectrally isolated in-gap edge modes, we identified an counterintuitive in-gap state that remains exponentially localized but whose probability peak can be shifted by varying only the imaginary gauge field, in contrast to a second mode that stays pinned to the boundary. We further showed that this steerable in-gap mode can be dynamically prepared from a bulk wave packet by applying weak on-site gain at a single fixed site, chosen where the corresponding Hermitian eigenstate attains its maximum probability density. With the gain site kept unchanged, varying the gauge field produces long-time states that remain exponentially localized but peak at different positions. These results provide a simple route to generating position-tunable, exponentially localized in-gap states in non-Hermitian quasiperiodic lattices.
@@ Acknowledgements.
X.L.Z. thanks Zhi Li and Peng Zou for helpful discussions. This work was supported by the Joint Fund of Natural Science Foundation of Shandong Province (Grant No. ZR2024LLZ004), the National Natural Science Foundation of China(Grant No. 12575010, No. 12005110), and the Natural Science Foundation of Shandong Province (Grant No. ZR2020QA078, No. ZR2023QB065, No. ZR2023MD064).